# Quantum-Limited Acoustoelectric Amplification in a Piezoelectric-2DEG Heterostructure

### v2

@@ V Clamping
Having determined the amplifier gain, we now shift our focus to deriving the clamping condition. Here, it is essential to account for the general time-evolution of the phonon mode through a beam-splitter Hamiltonian covering both the phonon and the 2DEG electron modes. Assuming a beam-splitter Hamiltonian consisting of NN pumped electrons each coupling with the phonon field through a time-varying coupling coefficient $g_{j}(t)$ (corresponding to the $j^{\textrm{th}}$ electron), the noise and the signal intensity feature the same gain per unit length (as shown in Appendix D), causing the signal-to-noise ratio to remain constant throughout the amplification process (analogously to a laser or maser). This common gain per unit length is given in terms of the coupling coefficient between a single electron and a phonon as follows:
@@
where $\tau$ is the time elapsed in a single amplifier segment (with a corresponding length of $v_{s}\tau$), $g=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$ is root-mean-squared (over the NN electrons) of the time-averaged amplitude (averaged over the period $\tau$) of the coupling coefficient, and $M=L/(v_{s}\tau)$ is the number of amplifier segments. We can conceptualize $\tau$ as the effective reset time for an electron after it emits a phonon (i.e., the time the electron takes to return to the original state). Given an electronic decay rate $\gamma/2$ for the initial and final states each, the overall reset time is given approximately as $\tau\approx 1/\gamma$.
Although the gain per unit length at low phonon intensity is proportional to $g^{2}N$, the clamping condition will be solely dependent on $g$ itself, as we will show later in this section. As such, we aim to dis-aggregate $g$ and NN by separately deriving $g$. We start by solving the interaction coefficient $g_{\mathrm{emit}}$ corresponding to phonon emission by pumped electrons. This coefficient acquires a time-dependence due to electronic decay and electron-phonon detuning, as shown in the following composite interaction Hamiltonian, where we sum over the interactions with all electrons in the region $S$ (see Fig. 3(a)):
@@ Appendix A Imaginary Part of Lindhard Chi(1) in High-Mobility/Low-Drift-Velocity Calculation
Here, we derive the imaginary part of $\chi^{(1)}_{L}$ in the high-mobility/low-drift-velocity limit. We start by integrating over the valid phase-space region as follows:
@@
where $\beta$ is a constant function of the 2DEG and piezoelectric material parameters:
@@
and $f^{(1)}_{k_{x}}$ corresponds to the closeness of the electron-phonon interaction to resonance and is thus a function of the electron transition frequency $\omega_{k_{x}-q,k_{x}}=\omega_{k_{x}-q}-\omega_{k_{x}}$:
@@
It is convenient to convert the phase-space integral in Eq. (75) to an integral over $\Delta\omega$. To this end, $dk_{x}$ is linear in $d(\Delta\omega)$, since any shift in the electron transition frequency $\omega_{k_{x},k_{x}-q}$ is proportional to the shift in $k_{x}$:
@@
Similarly, $k_{y,\mathrm{span}}$ can also be expressed in terms of the detuning $\Delta\omega$:
@@
Substituting the above expressions into Eq. (75), we calculate the imaginary part of $\chi^{(1)}_{L}$ in the following manner:
@@ Appendix B Imaginary Part of Lindhard Chi(1) in Low-Mobility Limit
We turn to deriving the imaginary part of $\chi^{(1)}_{L}$ in the low-mobility limit. In this regime, $f^{(1)}$ is written more fully as follows:
$f^{(1)}_{k_{x}^{\prime}}(-\omega_{0})=f_{1,k_{x}^{\prime}}^{(1)}(-\omega_{0})+f_{2,k_{x}^{\prime}}^{(1)}(-\omega_{0}),$ (83)
where $f_{1}$ and $f_{2}$ are the resonant (emissive) and counter-resonant (absorptive) terms, respectively:
@@
We solve for the net imaginary part of $\chi^{(1)}_{L}$ as follows by accounting for both the emission and absorption processes:
$\begin{split}&\textrm{Im}[\chi^{(1)}_{L}(-\omega_{0})]\\ &=\beta\bigg(\int_{S}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{1,k_{x}^{\prime}}(-\omega_{0})]+\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{2,k_{x}^{\prime}}(-\omega_{0})]\bigg)\\ &=\beta\bigg(\int_{S^{\prime}}dk_{x}^{\prime\prime}k_{y,\mathrm{span}}(-k_{x}^{\prime\prime})\textrm{Im}[f^{(1)}_{1,-k_{x}^{\prime\prime}}(-\omega_{0})]+\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{2,k_{x}^{\prime}}(-\omega_{0})]\bigg)\\ &=\beta\gamma\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\bigg(\frac{1}{(\omega_{-k_{x}^{\prime}+k_{d}-q,-k_{x}^{\prime}+k_{d}}+\omega_{0})^{2}+\gamma^{2}}-\frac{1}{(\omega_{k_{x}^{\prime}+k_{d}+q,k_{x}^{\prime}+k_{d}}-\omega_{0})^{2}+\gamma^{2}}\bigg)\\ &=\beta\gamma\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\bigg(\frac{1}{(\omega_{k_{x}^{\prime}-k_{d}+q,k_{x}^{\prime}-k_{d}}+\omega_{0})^{2}+\gamma^{2}}-\frac{1}{(\omega_{k_{x}^{\prime}+k_{d}+q,k_{x}^{\prime}+k_{d}}-\omega_{0})^{2}+\gamma^{2}}\bigg),\end{split}$ (86)
where in the second step, we established the coordinate $k_{x}^{\prime\prime}=-k_{x}^{\prime}$, whereas in the third step, we folded both of the integrals over $S^{\prime}$ into the single coordinate $k_{x}^{\prime}$ along with using the fact that $k_{y,\mathrm{span}}$ is even about $k_{x}^{\prime}=0$, and in the fourth step, we used the evenness of the electronic energy spectrum (i.e., the fact that $\omega_{k_{x,f},k_{x,i}}=\omega_{-k_{x,f},-k_{x,i}}$) to re-write the first term (representing the emission process).
@@ Appendix D Amplified Phonon Wave and Quantum Noise
Note that all NN electrons can potentially change to $|l\rangle_{k}$ state while emitting a phonon with an energy $\hbar\omega$. To be clear, we will divide the entire amplifier length into $M$ segments with identical lengths. NN is the number of “excited” electrons in each segment.
A quantum operator AA evolves in time as
@@
We start with the following beam-splitter Hamiltonian, consisting of a sum over NN electron-phonon interactions, each with a time-varying coupling coefficient $g_{j}(t)$:
@@
where $G_{0}=1+(g\tau)^{2}N$, and $g$ is defined as the square-root of the expectation value of $|g_{j}|^{2}$ over the NN electrons, i.e., $g=\left\langle|g_{j}|^{2}\right\rangle_{j}^{1/2}=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$. We also take the partial expected operator of the phonon field as:

### v3

@@ IV Calculating the Amplifier Gain
Finally, we briefly discuss the effect of temperature. As discussed in Sec. III, the variation of the Lindhard susceptibilities with temperature is mediated approximately fully through carrier density and mobility. This implies that the Mermin susceptibilities, and therefore the overall gain, are also similarly invariant in temperature if carrier density and mobility are given quantities, a conclusion borne out by the fact that the low-mobility gain equation calculated above matches room-temperature experimental results Wendt et al. (2025). It is also worth noting that since the initial phonon population in any given mode is set by the Bose-Einstein occupation number, the overall thermal energy gain for the system can be determined by performing a weighted sum of the single-mode gain derived above over all possible mode using the Bose-Einstein occupation numbers as the weighting factors.
@@ V Clamping
Having determined the amplifier gain, we now shift our focus to deriving the clamping condition. Here, it is essential to account for the general time-evolution of the phonon mode through a beam-splitter Hamiltonian covering both the phonon and the 2DEG electron modes. Assuming a beam-splitter Hamiltonian consisting of $N$ pumped electrons each coupling with the phonon field through a time-varying coupling coefficient $g_{j}(t)$ (corresponding to the $j^{\textrm{th}}$ electron), the noise and the signal intensity feature the same gain per unit length (as shown in Appendix D), causing the signal-to-noise ratio to remain constant throughout the amplification process (analogously to a laser or maser). This common gain per unit length is given in terms of the coupling coefficient between a single electron and a phonon as follows:
@@
where $\tau$ is the time elapsed in a single amplifier segment (with a corresponding length of $v_{s}\tau$), $g=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$ is root-mean-squared (over the $N$ electrons) of the time-averaged amplitude (averaged over the period $\tau$) of the coupling coefficient, and $M=L/(v_{s}\tau)$ is the number of amplifier segments. We can conceptualize $\tau$ as the effective reset time for an electron after it emits a phonon (i.e., the time the electron takes to return to the original state). Given an electronic decay rate $\gamma/2$ for the initial and final states each, the overall reset time is given approximately as $\tau\approx 1/\gamma$.
Although the gain per unit length at low phonon intensity is proportional to $g^{2}N$, the clamping condition will be solely dependent on $g$ itself, as we will show later in this section. As such, we aim to dis-aggregate $g$ and $N$ by separately deriving $g$. We start by solving the interaction coefficient $g_{\mathrm{emit}}$ corresponding to phonon emission by pumped electrons. This coefficient acquires a time-dependence due to electronic decay and electron-phonon detuning, as shown in the following composite interaction Hamiltonian, where we sum over the interactions with all electrons in the region $S$ (see Fig. 3(a)):
@@ Appendix A Imaginary Part of Lindhard Chi(1) in High-Mobility/Low-Drift-Velocity Calculation
Here, we derive the imaginary part of $\chi^{(1)}_{L}$ in the high-mobility/low-drift-velocity limit. We start by integrating over the valid phase-space region $S^{\prime}$ (see Fig. 3(b)) as follows:
@@
where $\beta$ is a constant function of the 2DEG and piezoelectric material parameters Chatterjee et al. (2024), proportional to the square of the dipole moment for an electron that has undergone interaction with the phonon field and inversely proportional to the 2DEG thickness:
@@
and $f^{(1)}_{k_{x}}$ corresponds to the closeness of the electron-phonon interaction to resonance (i.e., the inverse of the complex detuning) and is thus a function of the electron transition frequency $\omega_{k_{x}-q,k_{x}}=\omega_{k_{x}-q}-\omega_{k_{x}}$:
@@
It is convenient to convert the phase-space integral in Eq. (76) to an integral over $\Delta\omega$. We note that $dk_{x}$ is linear in $d(\Delta\omega)$, since any shift in the electron transition frequency $\omega_{k_{x},k_{x}-q}$ is proportional to the shift in $k_{x}$:
@@
In the high-mobility limit, the contributions to the integral will be dominated by the near-resonance region (see Fig. 4). There, the deviation of $k_{y,\mathrm{span}}$ from the value at resonance approximately varies linearly with $\Delta\omega$:
@@
Note that the $k_{y,\mathrm{span}}$ at resonance (i.e., $2m\omega^{\prime}/(\hbar k_{F})$) increases with the drift velocity in the low-drift-velocity limit, becoming positive for $v_{d}>v_{s}$. This matches the intuition laid out earlier in this work. Substituting the above expressions into Eq. (76), we calculate the imaginary part of $\chi^{(1)}_{L}$ in the following manner:
@@ Appendix B Imaginary Part of Lindhard Chi(1) in Low-Mobility Limit
We turn to deriving the imaginary part of $\chi^{(1)}_{L}$ in the low-mobility limit. In this regime, all possible electron-phonon interactions will fall within the near-resonance window. Therefore, $f^{(1)}$ needs to incorporate both resonant (emissive) and counter-resonant (absorptive) terms. We label these terms as $f_{1}$ and $f_{2}$, respectively:
@@
We solve for the net imaginary part of $\chi^{(1)}_{L}$ by integrating the resonant emissive terms and the counter-resonant absorptive terms over the regions $S$ and $S^{\prime}$, respectively (see Fig. 3):
$\begin{split}&\textrm{Im}[\chi^{(1)}_{L}(-\omega_{0})]\\ &=\beta\bigg(\int_{S}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{1,k_{x}^{\prime}}(-\omega_{0})]+\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{2,k_{x}^{\prime}}(-\omega_{0})]\bigg)\\ &=\beta\bigg(\int_{S^{\prime}}dk_{x}^{\prime\prime}k_{y,\mathrm{span}}(-k_{x}^{\prime\prime})\textrm{Im}[f^{(1)}_{1,-k_{x}^{\prime\prime}}(-\omega_{0})]+\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{2,k_{x}^{\prime}}(-\omega_{0})]\bigg)\\ &=\beta\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\textrm{Im}[f^{(1)}_{\mathrm{net},k_{x}^{\prime}}(-\omega_{0})]\\ &=\beta\gamma\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\bigg(\frac{1}{(\omega_{-k_{x}^{\prime}+k_{d}-q,-k_{x}^{\prime}+k_{d}}+\omega_{0})^{2}+\gamma^{2}}-\frac{1}{(\omega_{k_{x}^{\prime}+k_{d}+q,k_{x}^{\prime}+k_{d}}-\omega_{0})^{2}+\gamma^{2}}\bigg)\\ &=\beta\gamma\int_{S^{\prime}}dk_{x}^{\prime}k_{y,\mathrm{span}}(k_{x}^{\prime})\bigg(\frac{1}{(\omega_{k_{x}^{\prime}-k_{d}+q,k_{x}^{\prime}-k_{d}}+\omega_{0})^{2}+\gamma^{2}}-\frac{1}{(\omega_{k_{x}^{\prime}+k_{d}+q,k_{x}^{\prime}+k_{d}}-\omega_{0})^{2}+\gamma^{2}}\bigg),\end{split}$ (86)
where in the second step, we established the coordinate $k_{x}^{\prime\prime}=-k_{x}^{\prime}$, in the third step, we folded both of the integrals over $S^{\prime}$ into the single coordinate $k_{x}^{\prime}$ (using the fact that $k_{y,\mathrm{span}}$ is even about $k_{x}^{\prime}=0$) and defined the quantity $f_{\mathrm{net},k_{x}^{\prime}}^{(1)}=f_{1,-k_{x}^{\prime}}^{(1)}+f_{2,k_{x}^{\prime}}^{(1)}$, in the fourth step, we expressed the imaginary part of this quantity in terms of the electronic transition frequencies, and in the fifth step, we used the evenness of the electronic energy spectrum (i.e., the fact that $\omega_{k_{x,f},k_{x,i}}=\omega_{-k_{x,f},-k_{x,i}}$) to re-write the first term (representing the emission process).
@@ Appendix D Amplified Phonon Wave and Quantum Noise
Note that all $N$ electrons can potentially change to $|l\rangle_{k}$ state while emitting a phonon with an energy $\hbar\omega$. To be clear, we will divide the entire amplifier length into $M$ segments with identical lengths. $N$ is the number of “excited” electrons in each segment.
A quantum operator $A$ evolves in time as
@@
We start with the following beam-splitter Hamiltonian, consisting of a sum over $N$ electron-phonon interactions, each with a time-varying coupling coefficient $g_{j}(t)$:
@@
where $G_{0}=1+(g\tau)^{2}N$, and $g$ is defined as the square-root of the expectation value of $|g_{j}|^{2}$ over the $N$ electrons, i.e., $g=\left\langle|g_{j}|^{2}\right\rangle_{j}^{1/2}=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$. We also take the partial expected operator of the phonon field as: